Estimate the potential at 2cm from a proton and find the potential difference between two points that are 2cm and 4cm from a proton.


1
Expert's answer
2021-01-13T11:35:02-0500

Answer

Potential difference due to a point charge is given

"V=\frac{kq}{r}"

the potential at 2cm from a proton

"V=\frac{kq}{r}=\frac{9\times10^9\times1.6\times10^{-19}}{0.02}"

"=72\times10^{-9}V"

the potential at 4cm from a proton

"V'=\frac{kq}{r}=\frac{9\times10^9\times1.6\times10^{-19}}{0.04}"

"=36\times10^{-9}V"

the potential difference between two points that are 2cm and 4cm from a proton is given

"\Delta V=V-V'"

"=(72-36)\times10^{-9}V\\=36\times10^{-9}V"